### v1

# An Enhanced Projection Pursuit Tree Classifier with Visual Methods for Assessing Algorithmic Improvements

Abstract
This paper presents enhancements to the projection pursuit tree classifier and visual diagnostic methods for assessing their impact in high dimensions. The original algorithm uses linear combinations of variables in a tree structure where depth is constrained to be less than the number of classes—a limitation that proves too rigid for complex classification problems. Our extensions improve performance in multi-class settings with unequal variance-covariance structures and nonlinear class separations by allowing more splits and more flexible class groupings in the projection pursuit computation. Proposing algorithmic improvements is straightforward; demonstrating their actual utility is not. We therefore develop two visual diagnostic approaches to verify that the enhancements perform as intended. Using high-dimensional visualization techniques, we examine model fits on benchmark datasets to assess whether the algorithm behaves as theorized. An interactive web application enables users to explore the behavior of both the original and enhanced classifiers under controlled scenarios. The enhancements are implemented in the R package PPtreeExt.
1 Introduction
Loh (2014) provides an extensive overview of decades of research on classification and regression tree algorithms. Tree models are useful because they are simple, yet flexible, fast to compute, widely applicable and interpretable. A main drawback, though, most available methods use splits along single variables, which means that separations best made using a combination of variables are awkward for trees to fit. While there have been many suggestions it is a much harder problem to optimize on linear combinations, and thus there is less available successful software for oblique splits.
One available method and software is the Projection Pursuit Tree (PPtree) algorithm (Y. D. Lee et al. 2013). It optimizes a projection pursuit index, for example the LDA (E.-K. Lee et al. 2005) or PDA (E.-K. Lee and Cook 2010) index among others, based on class information to identify one-dimensional projections that best separate the groups at each node. The PPtree tree structure is simpler than classic methods like rpart (Therneau and Atkinson 2025), as it restricts the depth to G1G-1 ($G$ is the number of classes). At each node, PPtree separates the data into two groups by first deciding which classes to group together; the data are split into left and right branches based on two group means. This produces a compact and interpretable tree implemented in PPtreeViz R package (E.-K. Lee 2018).
One of the main advantages of PPtree is its ability to leverage the correlation between predictor variables to improve class separation. It has been shown to outperform even random forests (RF) in some scenarios. Projection pursuit addresses a known limitation of the original RF algorithm, in which oblique projections were allowed. While it sounds appealing, the algorithm is practically infeasible because it is based on randomly generated splits. Thus it is slow, and inefficient, for finding useful separating oblique splits.
The compact nature of PPtree means that it fits simply, and does not require pruning. The resulting one-dimensional projections can be used to make visualizations of group separations. However, there are limitations that reduce its effectiveness in certain contexts. Due the the way classes are grouped, the prediction boundaries are often too close to one group, resulting in unnecessarily high error rates. This happens because grouping of classes likely produces unequal variance-covariances, making the pooled variance-covariance used by LDA inadequate. Moreover, when classes have non-elliptical variance-covariance, or are split into multiple clusters, the rigidity of the (G(G-1)-node structure, will not be sufficiently flexible.
A primary objective of this paper is to describe modifications to the PPtree algorithm that address these limitations. The algorithm is changed in two ways: (1) improving prediction boundaries by modifying the choice of split points-through class subsetting; and (2) increasing flexibility by allowing multiple splits per group. The modified algorithm is tested on a variety of benchmark data sets. These data sets include well-separated classes, the presence of outliers, and nonlinear boundaries. The performance is compared with the original PPtree algorithm and existing methods. Visual comparisons of decision boundaries, along with error rates, are used to further understand the model fits. The ultimate goal is to provide better building blocks for a projection pursuit-based classification random forest algorithm (da Silva, Cook, and Lee 2021) implemented in PPforest R package (da Silva, Cook, and Lee 2025).
Proposing enhancements to algorithms is easy, and theoretically the modifications may seem sensible. In practice, they may not work as one envisions. Thus, an important part of developing algorithm modifications is providing validation that they are doing what is expected. This means going beyond predictive accuracy, to understanding the resulting models. A second objective of this paper is to provide visual diagnostics for the algorithm changes. This is two-fold: a shiny (Chang et al. 2025) app and tours to view the fits in high dimensions. The app is available in the PPtreeExt package (da Silva, Cook, and Lee 2026) along with the implementation of the modified algorithm. It allows users to generate 2D data sets with different types of separations and compare the model fits for the modified and the original algorithm. This was an essential tool in constructing the new algorithm, because it provides side-by-side visualization of prediction boundaries to identify data patterns in which the different methods perform better. The second visual diagnostic is to use tours available in the tourr package (Wickham et al. 2011) to examine the high-dimensional structure of data sets and the resulting model fits where the modified algorithm outperforms other classifiers. This provides an explanation for the performance comparison results, and intuition for why the modified algorithm performs better.
The paper is organized as follows: Section 2 provides background on the original PPtree algorithm, including its structure and limitations. Section 3 describes the extensions to the algorithm. Section 4 presents the interactive visualization tool developed for exploring and comparing the algorithms. Section 5 presents the comparative analysis of algorithm performance, and visual diagnostics for the comparisons in Section 6. Finally, Section 7 discusses the changes, results of the comparison and potential future directions.
2 Background to PPtree
Figure 1 illustrates the original PPtree algorithm for $G=3$. Starting from the optimal projection, the green group-having the most distant mean—is relabeled as $g_{1}^{*}$, while the orange and violet groups are combined and relabeled as $g_{2}^{*}$. A second projection vector is then computed using all observations, following the same procedure applied to these two new super-groups. A split point is determined using one of the predefined rules, and observations are assigned to branches by comparing their projected values to the cutoff.
The algorithm computes a split point using one of several methods and the observations are allocated to groups based on whether their projected values fall below or above the cutoff. Since $g_{1}^{*}$ contains only one of the original classes, the left branch becomes a terminal node. Because $g_{2}^{*}$ still includes two original classes, the procedure is recursively repeated on the right branch.
Figure 1: Illustration of the original PPtree algorithm for $G=3$.
There are eight different rules available for computing the split point on a given projection. Let $\bar{\mathbf{x}}_{g}$, $\mathbf{x}_{g}^{med}$, $s_{g}$, $IQR_{g}$, and $n_{g}$ denote the mean, median, standard deviation, interquartile range, and sample size of group $g$, respectively. Table LABEL:tbl-rules summarizes the formulas used to calculate the split point value, $c$, under each rule.
Table 1: Rules for computing the split value, $c$, between two groups, on a data projection, using means, weighted mean, medians, standard deviations, interquartile ranges.
Rule Name Formula
1 mean $c=\frac{1}{2}\bar{\mathbf{x}}_{1}+\frac{1}{2}\bar{\mathbf{x}}_{2}$
2 sample size weighted mean $c=\left(\frac{n_{2}}{n_{1}+n_{2}}\right)\bar{\mathbf{x}}_{1}+\left(\frac{n_{1}}{n_{1}+n_{2}}\right)\bar{\mathbf{x}}_{2}$
3 standard deviation weighted mean $c=\left(\frac{s_{2}}{s_{1}+s_{2}}\right)\bar{\mathbf{x}}_{1}+\left(\frac{s_{1}}{s_{1}+s_{2}}\right)\bar{\mathbf{x}}_{2}$
4 standard error weighted mean $c=\frac{s_{2}/\sqrt{n_{2}}}{s_{1}/\sqrt{n_{1}}+s_{2}/\sqrt{n_{2}}}\bar{\mathbf{x}}_{1}+\frac{s_{1}/\sqrt{n_{1}}}{s_{1}/\sqrt{n_{1}}+s_{2}/\sqrt{n_{2}}}\bar{\mathbf{x}}_{2}$
5 median $c=\frac{1}{2}med_{1}+\frac{1}{2}med_{2}$
6 sample size weighted median $c=\frac{n_{2}}{n_{1}+n_{2}}med_{1}+\frac{n_{1}}{n_{1}+n_{2}}med_{2}$
7 IQR weighted median $c=\frac{IQR_{2}}{IQR_{1}+IQR_{2}}med_{1}+\frac{IQR_{1}}{IQR_{1}+IQR_{2}}med_{2}$
8 sample size and IQR weighted median $c=\frac{IQR_{2}/\sqrt{n_{2}}}{IQR_{1}/\sqrt{n_{1}}+IQR_{2}/\sqrt{n_{2}}}med_{1}+\frac{IQR_{1}/\sqrt{n_{1}}}{IQR_{1}/\sqrt{n_{1}}+IQR_{2}/\sqrt{n_{2}}}med_{2}$
2.1 Algorithm
The PPtree algorithm (Y. D. Lee et al. 2013) follows a multi-step procedure to identify linear combinations of predictors that optimally separate classes in multi-class settings. Let $d_{n}=\{(\mathbf{x}i,y_{i})\}_{i=1}^{n}$ denote the dataset, where $\mathbf{x}_{i}$ is a $p$-dimensional vector of explanatory variables. The corresponding class label is $y_{i}\in\mathscr{G}$, with $\mathscr{G}=\{1,2,\ldots,G\}$, for $i=1,\ldots,n$. The construction of a PPtree classifier follows a four-step algorithm, described as follows:
PPtree Algorithm (4-step version)
Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
Output: A projection pursuit classification tree with at most G1G-1 internal nodes
  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates all classes in the current node.
  • Step 2: Binary Relabeling. On projected data, $z_{i}=\boldsymbol{\alpha}^{T}\mathbf{x}_{i}$, compute projected class means $\bar{z}_{1},\ldots,\bar{z}_{G}$. Identify the pair with the largest mean difference and compute the midpoint $mp$. Relabel the observations into two super-classes, $g_{1}^{*}$ and $g_{2}^{*}$, depending on whether $\bar{z}_{g}$ is smaller than the midpoint $mp$.
  • Step 3: Node Splitting. Optimize a new projection $\boldsymbol{\alpha}^{*}$ to separate $g_{1}^{*}$ and $g_{2}^{*}$ based on the relabeled classes. Define the split value $c$ on $z_{i}^{*}=\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{gi^{*}}$ using a predefined rule (e.g., mean or weighted mean). Split the node using the rule $z_{i}^{*}<c$.
  • Step 4: Recursive Partitioning. Repeat Steps 1–3 for each child node until every terminal node contains only one original class.
2.2 Less desirable aspects
The decision boundaries of a classifier reflect how the model partitions the feature space and adapt to class separations. Visualizing the decision boundaries provides valuable insights into model behavior and limitations. To illustrate the motivation for extending the PPtree algorithm, we present simple examples based on simulated datasets. Details on the simulation settings are provided in Section 4.
Figure 2 shows the decision boundaries produced by rpart (left) and PPtree (right). In this example, the class structure is best separated using linear combinations of the two variables. The rpart algorithm, which relies on axis-aligned splits, must approximate the ideal boundary using multiple steps. In contrast, PPtree generates an oblique split that more effectively captures the separation, although the resulting boundary between the orange and purple classes lies too close to one of them.
The first split isolates the violet class from the other two. A second partition then attempts to separate the green and orange groups. Although the group structure is similar across the three classes, the resulting decision boundaries are not parallel. This occurs because the initial split uses the combined information from the orange and green groups-treated as a super-group—to compute the projection and determine the splitting point.
Alternative rules for computing the split value (see Table LABEL:tbl-rules) have a substantial impact on the position and orientation of these decision boundaries.
Figure 2: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data. The boundaries generated by PPtree are oblique to the coordinate axes, capturing the linear association between the two variables.
Figure 3: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data. The orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node.
Figure 3 displays a simulated example that highlights the limited flexibility of the decision boundaries produced by PPtree. In this scenario, the orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node. In contrast, the more traditional rpart algorithm performs better in this case, as its recursive structure allows it to accommodate the relatively simple non-linear separation.
3 PPtree extensions
The algorithm has been extended in two main ways: by modifying the rules used to determine decision boundaries and by allowing multiple splits per group. Two new approaches are introduced for computing the optimal split value. To support multiple splits within a group, additional stopping rules must be incorporated into the algorithm.
3.1 Modification 1: Subsetting classes to produce better boundaries
The first modification targets the third step of the original algorithm. Rather than combining all remaining classes into a single super-class, only the two closest ones—identified by the smallest difference in their projected means—are used to compute the next projection and define the split point
$3^{*}$ Find an optimal one-dimensional projection $\alpha^{**}$, using a subset of $\{(\mathbf{x_{i}},y_{i}^{*})\}_{i=1}^{n}$ to separate the two class problem $g_{1}^{*}$ and $g_{2}^{*}$. To this second projection only information from the two closest groups, one from $g_{1}^{*}$ and the other from $g_{2}^{*}$ are used. The best separation of $g_{1}^{*}$ and $g_{2}^{*}$ is determined by $\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{g1^{*}}<c$; then $g_{1}^{*}$ is assigned to the left node else assign $g_{2}^{*}$ to the right node, where $\mathbf{\bar{x}}_{g1^{*}}$ is the mean of $g_{1}^{*}$.
Figure 4 illustrates the modified version of the algorithm for a three-class classification problem. After identifying the initial optimal projection, instead of combining two groups into a single super-group, the group furthest from the others is temporarily excluded. In this example, class $g_{3}$ is dropped from the current node. A new projection is then computed using only classes $g_{1}$ and $g_{2}$ (relabeled as $g_{1}^{*}$ and $g_{2}^{*}$ ), and one of the eight available rules is applied to determine the optimal split point. Subsequently, class $g_{3}$ is reintroduced, and a new projection is computed to best separate it from $g_{2}$, completing the tree construction.
Figure 6 shows the effect of this modification on the decision boundaries obtained from the simulated dataset presented in Figure 2.
Figure 4: Illustration of the algorithm with Modification 1 applied to a three-class problem. At each node, the second one-dimensional projection is computed using only the two closest groups to determine the best projection direction and split point.
3.2 Modification 2: Multiple splits per class based on entropy reduction
This modification introduces a new approach for selecting the split value, based on the impurity of the resulting groups. This approach resembles the impurity-based criteria used in the rpart algorithm, but it operates on projected data rather than individual features. Currently, the only implemented impurity measure is negative entropy:
$E(s)=-\sum_{j=1}^{G}p_{js}\log(p_{js})$ (1)
where $p_{js}$ is the proportion of observations from class $j$ in subset $s$, and $G$ is the number of classes. Higher values of $E(s)$ indicate greater class impurity. When all observations in the subset belong to the same class, the entropy reaches its minimum value of $E(s)=0$, indicating a pure node. To evaluate the quality of a candidate split, the impurity values of the resulting left and right subsets are combined using a weighted average of the entropie:
$E(s_{L},s_{R})=\frac{n_{L}}{n_{L}+n_{R}}E(s_{L})+\frac{n_{R}}{n_{L}+n_{R}}E(s_{R})$ (2)
where $n_{L},n_{R}$ are the number of observations in the left and right subsets, respectively. The combined impurity, $E(s_{L},s_{R})$ is computed for all possible split points-defined as the midpoints between consecutive ordered projected values-denoted $\mathscr{C}=\{c_{1},...,c_{n_{L}+n_{R}-1}\}$. The split point that minimizes the combined entropy is selected:
${\min_{c\in\mathscr{C}}}E_{c}(s_{L},s_{R})$
This represents a major departure from the original PPtree structure, aiming to enhance flexibility in modeling non-linear class boundaries by allowing multiple splits within a class. Instead of relying on summary statistics at the class level, this approach shifts the focus to local differences between individual observations. To support this flexibility, additional control mechanisms are required, particularly new stopping criteria. While this is conceptually similar to the rpart algorithm, a key distinction is that larger sample sizes per node are needed to reliably estimate projection coefficients.
Figure 5 illustrates this modified version of the algorithm, which follows the steps outlined below.
PPtree Algorithm with Entropy-Based Splitting
Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
Output: A projection pursuit classification tree with entropy-based node splitting
  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates the classes in the current node.
  • Step 2: Projection and Candidate Splits. On the projected data, $z_{i}=\boldsymbol{\alpha}^{T}\mathbf{x}_{i}$, define candidate splits as midpoints between consecutive $z_{i}$ values.
  • Step 3: Impurity Evaluation. For each candidate split $c$, divide the data into $s_{L}=\{z_{i}<c\}$ and $s_{R}=\{z_{i}\geq c\}$. Compute the class proportions $p_{js}$ for each subset and evaluate the entropy:
$E(s)=-\sum_{j=1}^{G}p_{js}\log(p_{js})$
Compute the total impurity as:
$E(s_{L},s_{R})=\frac{n_{L}}{n_{L}+n_{R}}E(s_{L})+\frac{n_{R}}{n_{L}+n_{R}}E(s_{R})$
  • Step 4: Best Split Selection. Select the split point $c^{*}$ that minimizes $E(s_{L},s_{R})$. Split the node using the rule $z_{i}<c^{*}$.
  • Step 5: Recursive Partitioning. Repeat Steps 1–4 recursively on each child node until some of the following stopping criteria are met:
    • All observations in the node belong to the same class;
    • The node size is smaller than a minimum threshold $n_{s}$;
    • The entropy reduction is smaller than a threshold $ent_{s}$.
Figure 5: Illustration of the algorithm with Modification 2 applied to a three-class problem. Projections of the data are computed at each node, and multiple splits per class are allowed. An impurity-based criterion, such as entropy, is used to determine when and how often to split observations within a node.
4 Algorithm comparison
To evaluate both the original algorithm and the proposed modifications, we developed an interactive Shiny app (Chang et al. 2025). The tool includes methods for simulating 2D classification problems, and displays the decision boundaries generated by different algorithms: CART, PPtree, and the proposed extensions (implemented in rpart, PPtreeViz, and PPtreeExt, respectively). The Shiny app is available at https://natyds.shinyapps.io/explorapp/ and is implemented in PPtreeExt.
The app contains three tabs, each tab offers control over data generation and algorithm parameters: Basic-Sim, Sim-Outlier, and MixSim. The first tab simulates well-separated classes, the second includes a cluster of outliers, and the third generates overlapping groups using the MixSim package (Melnykov, Chen, and Maitra 2012). Each tab offers control over data generation and algorithm parameters. The decision boundaries for rpart, PPtree, and the modified PPtree are displayed side by side for visual comparison.
Additional controls, such as the rule partition used in PPtree algorithm and the modified PPtree based on subsetting clases, can be incorporated into the application. These features allow users to examine how changes in rule criteria influence the resulting classification boundaries.
Figure 6 illustrates the use of the interactive application. The first modification (Subsetting clases) is applied to generate the projection, and Rule 1 is used to compute the split point. The modified PPtree achieves an error rate comparable to the original PPtree, with both substantially outperforming rpart. Additionally, the resulting decision boundary more accurately partitions the groups in a balanced manner.
Figure 6: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree based on subsetting clases applied to simulated well-separated classes. For PPtree, Rule 1 is used to determine the split points. The modified PPtree produces decision boundaries that are more symmetrically placed between the clusters.
Figure 7 presents a scenario in which Modification 2 (Multiple splits) yields a substantial improvement over the other algorithms. The use of multiple splits enables the small group of outliers from class 2 to be correctly classified.
Figure 7: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree—applied to data in which one class consists of two distinct clusters. The boundaries produced by the modified PPtree with multiple splits better capture the underlying structure of the data compared to the other methods.
Figure 8 compares the performance of the algorithm variants on data simulated from Gaussian mixtures. The user can adjust several parameters of the mixture model: $n$ is the sample size, $BarOmega$ is the desired average overlap between components, $MaxOmega$ specifies the maximum allowable overlap, and $K$ is the number of groups.
Figure 8: Decision boundaries produced by the three algorithms— rpart, PPtree, and the modified PPtree—on data simulated from a Gaussian mixture model.
The interactive application provides insight into how modifications to the algorithm affect model fit. It follows the principle of visualizing the model in the data space (Wickham, Cook, and Hofmann 2015). An additional and important use of the tool has been for debugging and evaluating the algorithmic behavior. By adjusting the simulated data and model parameters, it becomes evident that small changes in the data can result in substantial changes in the fitted model—particularly when projections are involved.
This sensitivity is likely to be even greater in higher-dimensional settings. However, in scenarios where class separation occurs along combinations of variables, the use of projected data at each node tends to yield decision boundaries that more closely align with the true data distribution.
5 Performance comparison
This section presents a benchmark data study to evaluate the predictive performance of the proposed method. We compare the extended PPtree models with several classification methods—CART, PPtree, Random Forest, PPforest, and SVM—using 94 benchmark datasets. Table LABEL:tbl-source-tab1 in the Appendix contain the url with the source data while Table LABEL:tbl-data-tab2 summarize the main characteristics of each dataset summarizing the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes, biggest proportion size difference between classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
For each dataset, two-thirds of the observations are randomly selected for training, and the remaining one-third is used for testing and computing the prediction error. This process is repeated 200 times.
Figure 9: Average test error from 200 training/test splits of benchmark data for four models: rpart, PPtree PPtreeExt modifications 1 and 2, where the PPtree extension performs better than rpart and the original PPtree.
Moreover, in 18 datasets, the modified PPtree outperforms CART, as illustrated in Figure 9, and in only 3 cases it achieves the lowest error among all selected methods, including ensemble-based classifiers. Table LABEL:tbl-datatbl presents an overview of these datasets, which is a subset of Table LABEL:tbl-data-tab2 from the Appendix.
Table 2: Overview of benchmark datasets where the proposed tree-based extensions outperform the selected tree methods. The table summarizes the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
dataset Cases Predictors Groups Imbalance Correlation
data20-3 8143 5 2 0.58 0.45
data27-a11 1747 18 5 0.38 0.24
data27-a31 1834 18 5 0.28 0.27
data27-b31 1424 18 5 0.21 0.24
data27-c11 1111 18 5 0.14 0.27
data27-c31 1448 18 5 0.17 0.27
data37-5 17560 53 5 0.21 0.16
data38 182 12 2 0.43 0.21
data63 6598 166 2 0.69 0.27
data34 1372 4 2 0.11 0.43
data49-2 606 100 2 0.03 1.00
data49-3 606 100 2 0.01 0.99
data22-1 258 5 4 0.25 0.12
data49-4 606 100 2 0.01 0.99
data56-1 7494 16 10 0.01 0.27
data67 20000 16 26 0.00 0.18
data69-1 5000 21 3 0.01 0.30
data69-2 5000 40 3 0.01 0.09
6 Explanation of performance
We would expect that the modified PPtree algorithm outperforms
  • rpart when the best separation between groups is in oblique directions
  • and the original PPtree when class variances are different, or when there are multiple classes.
This can be investigated visually using a tour. There are several strategies for approaching this:
  1. 1.
    On the data
    1. a.
      Work on two groups at a time, if there are more than 2
    2. b.
      Use a guided tour on the data to find the combinations of the variables where the groups are separated
  2. 2.
    Predict a grid of values that covers the data domain, for the methods interested in comparing, PPtreeExt, rpart and PPtree
    1. a.
      subset to the points near the border generated by the classifier.
    2. b.
      record a sequence of projection bases that shows the boundary created by PPtree
    3. c.
      use the same projection sequence to view the boundaries for the three different methods, possibly use a slice tour to focus on the boundary near the center of the data domain.
Figure 10 show projections of the data20-3 illustrating why PPtreeExt performs better. Group 0 is split on either side of group 1, which means that the original PPtree algorithm would not be able to capture both clusters. We can also see that the separation is mostly on V3 but small amounts of other variables help to improve the separation. This means that rpart will not work quite as well because it will focus on splitting V3.
(a) projection 1
(b) projection 2
Figure 10: Two projections from a grand tour of data20-3, from which we can see why the PPtreeExt method is the most effective. The group 0 is split into two clusters on either side of group 1, which means that the original algorithm is not optimal. Secondly, the separation between the two groups is in a combination of variables.
Figure 11 shows some plots illustrating why PPtreeExt outperforms the other methods of the data49-2 data. From Figure 9 we can see that the original PPtree (with LDA) algorithm performs much better than a tree, and the PPtreeExt performs much better again. This data has 100 predictors and 2 classes (Table LABEL:tbl-datatbl), which makes visualising more challenging that data20-3. A useful approach is to pre-process using principal component analysis (PCA) on the predictors, to reduce the number of variables to visualise, while maintaining class structure. Plot (a) shows PC 1 vs PC2, with points coloured by the two classes. PC 1 explains 99.8% of the variability in the predictors but it explains nothing of the class structure. This is not uncommon for PCA, where often interesting structure can be found in the smaller components. Plot (b) shows PC 2 vs PC 3, and here we see interesting class differences. Each class appears to have three linear pieces pointing in different directions. This supports why PPtreeExt might work well, because in the combination of variables, the classes are nearly separable with multiple linear cuts. It’s not feasible to continue working through principal components to tease apart the cluster structure relating to classes, so a good next step is to do projection pursuit. The optimal projection is shown in plot (c). It can be seen the two classes are mostly separable by a linear combination of variables, but there is no gap between the clusters. The linear combination with so many variables is what causes some difficulties for a simple tree algorithm. The interesting shapes of the clusters for each class, which appear to be multiple linear pieces, means that the PPtreeExt can find better separations than PPtree.
(a) PC1 vs PC2
(b) PC2 vs PC3
(c) optimal projection
Figure 11: Examining why data49-2 is best classified using PPtreeExt. Because there are 100 predictors, for visualisation it is useful to reduce the dimension using principal component analysis, and then examine the separation between groups. PC1 explains 99.8% of the overall variability among predictors, but it is not useful for separating the classes (a). PC2 and PC3 (b) have an interesting difference between the two classes which shows why multiple linear splits might be useful. A guided projection pursuit on the first 26 principal components (c) shows that these classes could be separated with a linear combination. This is why the original PPtree classifier does better than the tree algorithm and also why the extension improves the performance.
Figure 12 has plots illustrating why the PPtreeExt performs better on data69-1. From Figure 9 we see that all of the PPtree versions perform better than an rpart fit. This data has 21 predictors and 3 classes (Table LABEL:tbl-datatbl). This is sufficiently small to examine using a tour, and we initially use a guided tour with the LDA index. This produced the projection shown in plot (a) where we see the primary difference between the three classes. These are each similar size, but oriented in such a way that in a 2D projection they form a triangle. Plots (b) and (c) show the boundaries using a slice tour with the same projection as used in (a) for PPtreeExt and rpart, respectively. The boundary points are generated by simulating a random sample of observations between (-3, 3) in each dimension. The slice cuts through to centre of the box. We can see that PPtreeExt essentially fits the data shape neatly but the rpart boundary is very messy. This data needs oblique cuts, which PPtreeExt provides, and the modified algorithm handles the variance differences better than the original.
(a) data
(b) PPtreeExt
(c) rpart
Figure 12: Examining why data49-2 is best classified using PPtreeExt: (a) best projection from a guided tour shows that the data has three similar linear pieces oriented to form a triangle, (b) boundary produced by the PPtreeExt fitted model and (c) boundary produced by rpart fitted model.
7 Discussion
This article has presented two modifications to the PPtree algorithm for classification problems implemented in the R package PPtreeExt and evaluated using a purpose-built interactive web application and real data. The first modification involves subsetting classes when calculating the projection direction at each node, while the second allows for multiple splits per class based on individual entropy evaluations between candidate splits. These changes result in a more flexible classification method that leverages linear combinations of variables. The web application also demonstrates how easily diagnostic tools can be constructed to explore and understand the behavior of new algorithms.
The benchmark data study showed that PPtreeExt predictive performance is better than, CART and PPtree in 18 data out of 94. The modified PPtree algorithm outperforms CART in cases when the best separation between groups is in oblique directions and outperforms the original PPtree when class variances are different, or when there are multiple classes.
Interpretability is increasingly important, because predictions from algorithms may need to be defended in practice. We have shown how to examine the data and the fitted model to understand why the PPtree extension performs better than other tree methods on some data sets. This approach is generally useful for assessing and diagnosing model fits.
The modified tree algorithm can be incorporated into a projection pursuit forest classifier (da Silva, Cook, and Lee 2025). Given its implicit variable selection mechanism, the ensemble is expected to handle nuisance variables more effectively than a single tree. Additionally, the use of bagging should yield improved and more robust performance in problems with nonlinear decision boundaries. This can be accomplished using the current PPforest package.
8 Appendix
Table 3: Source data information
Name URL
data12 https://archive.ics.uci.edu/dataset/518/speaker+accent+recognition
data13-1 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
data13-2 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
data20-1 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data20-2 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data20-3 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data21 https://archive.ics.uci.edu/dataset/537/cervical+cancer+behavior+risk
data22-1 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data22-2 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data26 https://archive.ics.uci.edu/dataset/308/gas+sensor+array+under+flow+modulation
data27-a11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a21 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data29 https://archive.ics.uci.edu/dataset/282/lsvt+voice+rehabilitation
data32 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data34 https://archive.ics.uci.edu/dataset/267/banknote+authentication
data36 https://archive.ics.uci.edu/dataset/236/seeds
data37-1 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-2 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-3 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-4 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-5 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-8 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-14 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data38 https://archive.ics.uci.edu/dataset/230/planning+relax
data39-1 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
Name URL
data39-2 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-3 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-4 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-5 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-6 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-7 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-8 https://archive.ics.uci.edu/dataset/225/gas+sensor+array+drift+dataset
data39-9 https://archive.ics.uci.edu/dataset/226/gas+sensor+array+drift+dataset
data41-1 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data41-2 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data41-3 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data42 https://archive.ics.uci.edu/dataset/192/breast+tissue
data44 https://archive.ics.uci.edu/dataset/181/libras+movement
data49-2 https://archive.ics.uci.edu/dataset/167/hill+valley
data49-3 https://archive.ics.uci.edu/dataset/168/hill+valley
data49-4 https://archive.ics.uci.edu/dataset/169/hill+valley
data50 https://archive.ics.uci.edu/dataset/161/mammographic+mass
data52-1 https://archive.ics.uci.edu/dataset/96/spectf+heart
data52-2 https://archive.ics.uci.edu/dataset/96/spectf+heart
data53 https://archive.ics.uci.edu/dataset/139/synthetic+control+chart+time+series
data55-1 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
data55-2 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
data56-1 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
data56-2 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
data57 https://archive.ics.uci.edu/dataset/39/ecoli
data61-1 https://archive.ics.uci.edu/dataset/54/isolet
data62 https://archive.ics.uci.edu/dataset/74/musk+version+1
data63 https://archive.ics.uci.edu/dataset/75/musk+version+2
data64-1 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
data64-2 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
data65 https://archive.ics.uci.edu/dataset/15/breast+cancer+wisconsin+original
data66 https://archive.ics.uci.edu/dataset/62/lung+cancer
data67 https://archive.ics.uci.edu/dataset/59/letter+recognition
data68-1 https://archive.ics.uci.edu/dataset/50/image+segmentation
data68-2 https://archive.ics.uci.edu/dataset/50/image+segmentation
data69-1 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
data69-2 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
data70 https://archive.ics.uci.edu/dataset/42/glass+identification
data71-1 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
Name URL
data71-2 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
data72-1 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-2 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-3 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-4 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-5 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-6 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-7 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data73 https://archive.ics.uci.edu/dataset/151/connectionist+bench+sonar+mines+vs+rocks
data74 https://archive.ics.uci.edu/dataset/152/connectionist+bench+vowel+recognition+deterding+data
NCI60 https://github.com/natydasilva/PPforest/blob/master/data/NCI60.rda
crab https://github.com/natydasilva/PPforest/blob/master/data/crab.rda
fishcatch https://github.com/natydasilva/PPforest/blob/master/data/fishcatch.rda
glass https://github.com/natydasilva/PPforest/blob/master/data/glass.rda
image https://github.com/natydasilva/PPforest/blob/master/data/image.rda
leukemia https://github.com/natydasilva/PPforest/blob/master/data/image.rda
lymphoma https://github.com/natydasilva/PPforest/blob/master/data/lymphoma.rda
olive https://github.com/natydasilva/PPforest/blob/master/data/olive.rda
parkinson https://github.com/natydasilva/PPforest/blob/master/data/parkinson.rda
wine https://github.com/natydasilva/PPforest/blob/master/data/wine.rda
Table 6: Overview of benchmark datasets
dataset Cases Predictors Groups Imbalance Correlation
data12 329 12 6 0.41 0.38
data13-1 122 13 2 0.03 0.41
data13-2 121 13 2 0.29 0.37
data20-1 2665 5 2 0.27 0.81
data20-2 9752 5 2 0.58 0.29
data20-3 8143 5 2 0.58 0.45
data21 72 19 2 0.42 0.22
data22-1 258 5 4 0.25 0.12
data22-2 145 5 4 0.14 0.18
data26 58 435 4 0.21 0.64
data27-a11 1747 18 5 0.38 0.24
data27-a21 1264 18 5 0.35 0.22
data27-a31 1834 18 5 0.28 0.27
data27-a32 1830 32 5 0.28 0.13
data27-b11 1072 18 5 0.31 0.27
data27-b12 1069 32 5 0.32 0.13
data27-b31 1424 18 5 0.21 0.24
data27-b32 1420 32 5 0.21 0.11
data27-c11 1111 18 5 0.14 0.27
data27-c12 1107 32 5 0.13 0.11
data27-c31 1448 18 5 0.17 0.27
data27-c32 1444 32 5 0.17 0.11
data29 126 310 2 0.33 0.28
data32 403 6 4 0.20 0.11
data34 1372 4 2 0.11 0.43
data36 210 7 3 0.00 0.61
data37-1 29101 53 5 0.23 0.17
data37-14 769 53 2 0.52 0.16
data37-2 14132 53 3 0.62 0.18
data37-3 11835 53 4 0.34 0.16
data37-4 16357 53 5 0.29 0.15
data37-5 17560 53 5 0.21 0.16
dataset Cases Predictors Groups Imbalance Correlation
data37-8 22990 53 9 0.27 0.17
data38 182 12 2 0.43 0.21
data39-1 445 128 6 0.15 0.55
data39-2 1239 128 5 0.35 0.40
data39-3 1586 128 5 0.17 0.56
data39-4 161 128 5 0.32 0.43
data39-5 197 128 5 0.22 0.57
data39-6 2300 128 6 0.25 0.61
data39-7 3613 128 6 0.11 0.68
data39-8 294 128 6 0.43 0.57
data39-9 470 128 6 0.10 0.42
data41-1 5456 2 4 0.34 0.03
data41-2 5456 4 4 0.34 0.15
data41-3 5456 24 4 0.34 0.17
data42 106 9 6 0.08 0.51
data44 360 90 15 0.00 0.27
data49-2 606 100 2 0.03 1.00
data49-3 606 100 2 0.01 0.99
data49-4 606 100 2 0.01 0.99
data50 830 5 2 0.03 0.23
data52-1 80 44 2 0.00 0.25
data52-2 187 44 2 0.84 0.29
data53 600 60 6 0.00 0.46
data55-1 3823 64 10 0.00 0.12
data55-2 1797 64 10 0.01 0.12
data56-1 7494 16 10 0.01 0.27
data56-2 3498 16 10 0.01 0.28
data57 327 7 5 0.38 0.23
data61-1 6238 617 26 0.00 0.17
data62 476 166 2 0.13 0.26
data63 6598 166 2 0.69 0.27
data64-1 4435 36 6 0.15 0.49
data64-2 2000 36 6 0.13 0.50
data65 683 9 2 0.30 0.60
data66 27 56 3 0.07 0.20
dataset Cases Predictors Groups Imbalance Correlation
data67 20000 16 26 0.00 0.18
data68-1 210 19 7 0.00 0.30
data68-2 2100 19 7 0.00 0.28
data69-1 5000 21 3 0.01 0.30
data69-2 5000 40 3 0.01 0.09
data70 214 9 6 0.31 0.23
data71-1 43500 9 7 0.78 0.19
data71-2 14494 9 5 0.79 0.20
data72-1 94 18 4 0.09 0.39
data72-2 94 18 4 0.05 0.40
data72-3 94 18 4 0.10 0.42
data72-4 94 18 4 0.12 0.42
data72-5 94 18 4 0.10 0.49
data72-6 94 18 4 0.06 0.40
data72-7 94 18 4 0.06 0.44
data73 208 60 2 0.07 0.23
data74 990 12 11 0.00 0.22
NCI60 56 31 7 0.05 0.56
crab 200 6 4 0.00 0.95
fishcatch 159 7 7 0.31 0.46
glass 214 10 6 0.31 0.23
image 2310 19 7 0.00 0.28
leukemia 72 41 3 0.40 0.44
lymphoma 80 51 3 0.41 0.75
olive 572 9 9 0.32 0.35
parkinson 195 23 2 0.51 0.50
wine 178 14 3 0.13 0.30

### v2

@@
 An interactive web application enables users to explore the behavior of both the original and enhanced classifiers under controlled scenarios.
-The enhancements are implemented in the R package PPtreeExt.
+The enhancements are implemented in the R package PPtreeExt, which is available on CRAN.
+Natalia da Silva1   Dianne Cook2   Eun-Kyung Lee3
+1Instituto de Estadística (IESTA), Universidad de la República, Montevideo, Uruguay
+2Department of Econometrics and Business Statistics, Monash University, Melbourne, Australia
+3Department of Statistics, Ewha Womans University, Seoul, Republic of Korea
+Keywords: classification, heterogeneous variance, machine learning, projection pursuit, separability, tree classifier
 1 Introduction
 Loh (2014) provides an extensive overview of decades of research on classification and regression tree algorithms.
-Tree models are useful because they are simple, yet flexible, fast to compute, widely applicable and interpretable.
-A main drawback, though, most available methods use splits along single variables, which means that separations best made using a combination of variables are awkward for trees to fit.
-While there have been many suggestions it is a much harder problem to optimize on linear combinations, and thus there is less available successful software for oblique splits.
-One available method and software is the Projection Pursuit Tree (PPtree) algorithm (Y.
-D.
-Lee et al.
+Tree models are useful because they are simple, yet flexible, fast to compute, widely applicable, and interpretable.
+A main drawback, though, is that most available methods use splits along single variables, making separations on a combination of variables harder to capture.
+The reason is that it is a much harder problem to optimize the search for linear combinations, and thus, there is a lack of useful software that provides oblique splits.
+One available method and software is the projection pursuit tree (PPtree) (Lee et al.
 2013).
-It optimizes a projection pursuit index, for example the LDA (E.-K.
-Lee et al.
-2005) or PDA (E.-K.
-Lee and Cook 2010) index among others, based on class information to identify one-dimensional projections that best separate the groups at each node.
-The PPtree tree structure is simpler than classic methods like rpart (Therneau and Atkinson 2025), as it restricts the depth to G1G-1 ($G$ is the number of classes).
+It optimizes a projection pursuit (PP) index, for example, the LDA (Lee et al.
+2005) or PDA (Lee and Cook 2010) index among others, based on class information to identify one-dimensional projections that best separate the groups at each node.
+The PPtree tree structure is simpler than classic methods like rpart (Therneau and Atkinson 2025), as it restricts the tree depth to G1G-1 ($G$ is the number of classes).
 At each node, PPtree separates the data into two groups by first deciding which classes to group together; the data are split into left and right branches based on two group means.
-This produces a compact and interpretable tree implemented in PPtreeViz R package (E.-K.
-Lee 2018).
+This produces a compact and interpretable tree implemented in the PPtreeViz R package (Lee 2018).
 One of the main advantages of PPtree is its ability to leverage the correlation between predictor variables to improve class separation.
-It has been shown to outperform even random forests (RF) in some scenarios.
-Projection pursuit addresses a known limitation of the original RF algorithm, in which oblique projections were allowed.
-While it sounds appealing, the algorithm is practically infeasible because it is based on randomly generated splits.
-Thus it is slow, and inefficient, for finding useful separating oblique splits.
-The compact nature of PPtree means that it fits simply, and does not require pruning.
+In some scenarios, it has been shown to outperform other classifiers, including random forests (RF) (Breiman 2001).
+PP addresses a known limitation of the original RF algorithm.
+It included oblique projections, but because it is based on randomly generated splits was practically infeasible.
+It is an inefficient algorithm for finding useful oblique splits.
+The compact nature of PPtree means that it fits simply and does not require pruning.
 The resulting one-dimensional projections can be used to make visualizations of group separations.
-However, there are limitations that reduce its effectiveness in certain contexts.
-Due the the way classes are grouped, the prediction boundaries are often too close to one group, resulting in unnecessarily high error rates.
+However, there are limitations that reduce its effectiveness.
+In multiclass problems, due to the way classes are grouped, the prediction boundaries are often too close to one group, resulting in unnecessarily high error rates.
 This happens because grouping of classes likely produces unequal variance-covariances, making the pooled variance-covariance used by LDA inadequate.
-Moreover, when classes have non-elliptical variance-covariance, or are split into multiple clusters, the rigidity of the (G(G-1)-node structure, will not be sufficiently flexible.
-A primary objective of this paper is to describe modifications to the PPtree algorithm that address these limitations.
-The algorithm is changed in two ways: (1) improving prediction boundaries by modifying the choice of split points-through class subsetting; and (2) increasing flexibility by allowing multiple splits per group.
-The modified algorithm is tested on a variety of benchmark data sets.
-These data sets include well-separated classes, the presence of outliers, and nonlinear boundaries.
-The performance is compared with the original PPtree algorithm and existing methods.
-Visual comparisons of decision boundaries, along with error rates, are used to further understand the model fits.
-The ultimate goal is to provide better building blocks for a projection pursuit-based classification random forest algorithm (da Silva, Cook, and Lee 2021) implemented in PPforest R package (da Silva, Cook, and Lee 2025).
-Proposing enhancements to algorithms is easy, and theoretically the modifications may seem sensible.
+Secondly, when classes have non-elliptical variance-covariance, or have multiple clusters per class, the rigidity of the (G(G-1)-node structure is not sufficiently flexible.
+The algorithm is modified to address these limitations in two ways: (1) improving prediction boundaries by modifying the choice of split points through class subsetting; and (2) increasing flexibility by allowing multiple splits per group.
+Proposing enhancements to algorithms is easy, and theoretically, the modifications may seem sensible.
 In practice, they may not work as one envisions.
-Thus, an important part of developing algorithm modifications is providing validation that they are doing what is expected.
-This means going beyond predictive accuracy, to understanding the resulting models.
+Thus, an important part of developing algorithm modifications is providing validation that they are doing what is expected, or not.
+This means going beyond predictive accuracy to understanding the resulting models.
 A second objective of this paper is to provide visual diagnostics for the algorithm changes.
@@
 2025) app and tours to view the fits in high dimensions.
-The app is available in the PPtreeExt package (da Silva, Cook, and Lee 2026) along with the implementation of the modified algorithm.
+The app is available in the PPtreeExt package (da Silva et al.
+2026) along with the implementation of the modified algorithm.
 It allows users to generate 2D data sets with different types of separations and compare the model fits for the modified and the original algorithm.
-This was an essential tool in constructing the new algorithm, because it provides side-by-side visualization of prediction boundaries to identify data patterns in which the different methods perform better.
+This was an essential tool in constructing the new algorithm, because it provides a side-by-side visualization of prediction boundaries to identify data patterns in which the different methods perform better.
 The second visual diagnostic is to use tours available in the tourr package (Wickham et al.
-2011) to examine the high-dimensional structure of data sets and the resulting model fits where the modified algorithm outperforms other classifiers.
-This provides an explanation for the performance comparison results, and intuition for why the modified algorithm performs better.
+2011) to examine the high-dimensional structure of data sets and the resulting model fits, where the modified algorithm outperforms other classifiers.
+This explains the performance comparison results and intuition for why the modified algorithm performs better.
 The paper is organized as follows: Section 2 provides background on the original PPtree algorithm, including its structure and limitations.
 Section 3 describes the extensions to the algorithm.
-Section 4 presents the interactive visualization tool developed for exploring and comparing the algorithms.
-Section 5 presents the comparative analysis of algorithm performance, and visual diagnostics for the comparisons in Section 6.
-Finally, Section 7 discusses the changes, results of the comparison and potential future directions.
+Section 4 contains the examples on visually and numerically digesting the enhanced algorithm’s performance.
+Finally, Section 5 discusses the changes, results of the comparison, and potential future directions.
 2 Background to PPtree
 Figure 1 illustrates the original PPtree algorithm for $G=3$.
-Starting from the optimal projection, the green group-having the most distant mean—is relabeled as $g_{1}^{*}$, while the orange and violet groups are combined and relabeled as $g_{2}^{*}$.
+Starting from the optimal projection, the left-side, most distant, cluster (green circles) is relabeled as $g_{1}^{*}$, while the other two clusters are combined and relabeled as $g_{2}^{*}$.
 A second projection vector is then computed using all observations, following the same procedure applied to these two new super-groups.
 A split point is determined using one of the predefined rules, and observations are assigned to branches by comparing their projected values to the cutoff.
-The algorithm computes a split point using one of several methods and the observations are allocated to groups based on whether their projected values fall below or above the cutoff.
+The algorithm computes a split point using one of several methods, and the observations are allocated to groups based on whether their projected values fall below or above the cutoff.
 Since $g_{1}^{*}$ contains only one of the original classes, the left branch becomes a terminal node.
@@
 Figure 1: Illustration of the original PPtree algorithm for $G=3$.
+In step 1, $P$ optimization decides on the best linear projection.
+Step 2 combines two groups to make a two-group classification.
+The $P$ optimization is computed for two group problem, and an appropriate split is decided.
+This is repeated for the remaining two groups to get the final model containing two splits.
+The locations where modifications will be made are circled.
 There are eight different rules available for computing the split point on a given projection.
@@
 Table LABEL:tbl-rules summarizes the formulas used to calculate the split point value, $c$, under each rule.
-Table 1: Rules for computing the split value, $c$, between two groups, on a data projection, using means, weighted mean, medians, standard deviations, interquartile ranges.
+Table 1: Rules for computing the split value, $c$, between two groups, on a data projection, using means, weighted mean, medians, standard deviations, and interquartile ranges.
 Rule Name Formula
@@
 2.1 Algorithm
-The PPtree algorithm (Y.
-D.
-Lee et al.
+The PPtree algorithm (Lee et al.
 2013) follows a multi-step procedure to identify linear combinations of predictors that optimally separate classes in multi-class settings.
@@
 The corresponding class label is $y_{i}\in\mathscr{G}$, with $\mathscr{G}=\{1,2,\ldots,G\}$, for $i=1,\ldots,n$.
-The construction of a PPtree classifier follows a four-step algorithm, described as follows:
-PPtree Algorithm (4-step version)
+The construction of a PPtree classifier follows a four-step algorithm, as follows:
 Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
 Output: A projection pursuit classification tree with at most G1G-1 internal nodes
-  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates all classes in the current node.
+  • Step 1: Class Separation. Optimize a PP index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates all classes in the current node.
   • Step 2: Binary Relabeling.
@@
 2.2 Less desirable aspects
-The decision boundaries of a classifier reflect how the model partitions the feature space and adapt to class separations.
+The decision boundaries of a classifier reflect how the model partitions the feature space and adapts to class separations.
 Visualizing the decision boundaries provides valuable insights into model behavior and limitations.
-To illustrate the motivation for extending the PPtree algorithm, we present simple examples based on simulated datasets.
-Details on the simulation settings are provided in Section 4.
-Figure 2 shows the decision boundaries produced by rpart (left) and PPtree (right).
-In this example, the class structure is best separated using linear combinations of the two variables.
-The rpart algorithm, which relies on axis-aligned splits, must approximate the ideal boundary using multiple steps.
-In contrast, PPtree generates an oblique split that more effectively captures the separation, although the resulting boundary between the orange and purple classes lies too close to one of them.
-The first split isolates the violet class from the other two.
-A second partition then attempts to separate the green and orange groups.
+To illustrate the PPtree algorithm relative to the classical rpart and the motivation for improving PPtree, we use two simulated three-class 2D data sets.
+Data set LC has separations on linear combinations of variables (Figure 2), and data set MC has multiple clusters in a class (Figure 3).
+The three classes are indicated by squares (S, violet), circles (C, green), and triangles (T, orange).
+Figure 2 shows the decision boundaries produced by rpart (left) and PPtree (right), and percentage error, for the LC data.
+The class clusters are best separated using linear combinations of the two variables.
+The rpart algorithm, which relies on axis-aligned splits, must approximate the ideal boundary using rectangular steps.
+Even with many steps, it cannot be as efficient as the oblique splits provided by PPtree.
+The two steps of PPtree are to split the S cluster (violet) first, and then split C (green) and T (orange).
 Although the group structure is similar across the three classes, the resulting decision boundaries are not parallel.
-This occurs because the initial split uses the combined information from the orange and green groups-treated as a super-group—to compute the projection and determine the splitting point.
-Alternative rules for computing the split value (see Table LABEL:tbl-rules) have a substantial impact on the position and orientation of these decision boundaries.
-Figure 2: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data.
-The boundaries generated by PPtree are oblique to the coordinate axes, capturing the linear association between the two variables.
-Figure 3: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data.
-The orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node.
-Figure 3 displays a simulated example that highlights the limited flexibility of the decision boundaries produced by PPtree.
-In this scenario, the orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node.
+This occurs because the initial split uses the combined information from the C and T clusters - treated as a super class - to compute the projection and determine the splitting point.
+This also results in the boundary between T and S being too close to T.
+Alternative rules for computing the split value (see Table LABEL:tbl-rules) can impact the position and orientation of these decision boundaries, but in this case does not fix the deficiency.
+Figure 2: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on the LC data.
+The boundaries generated by PPtree are oblique to the coordinate axes, appropriately capturing the separations as linear combinations of the two variables.
+Even with many splits, the rpart algorithm cannot elegantly classify this data.
+Figure 3: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on the MC data.
+The T class (orange triangles) cannot be separated using a single linear partition, and PPtree fails to model it accurately because each class must be assigned to a single terminal node.
+This illustrates the need for allowing more splits in PPtree to capture unusually shaped clusters and separations involving combinations of variables.
+Figure 3 compares the rpart and PPtree decision boundaries for the data set MC, where class T has two separated clusters.
+It highlights the limited flexibility of the decision boundaries produced by PPtree.
+In this scenario, the T cannot be separated using a single linear partition, and PPtree fails to capture it because the tree depth is limited to a single terminal node for each class.
 In contrast, the more traditional rpart algorithm performs better in this case, as its recursive structure allows it to accommodate the relatively simple non-linear separation.
+However, rpart fails to capture the separation between classes as linear combinations.
+This shows why more splits should be allowed for the PPtree algorithm.
 3 PPtree extensions
-The algorithm has been extended in two main ways: by modifying the rules used to determine decision boundaries and by allowing multiple splits per group.
-Two new approaches are introduced for computing the optimal split value.
-To support multiple splits within a group, additional stopping rules must be incorporated into the algorithm.
+The algorithm has been extended in two main ways, as indicated by the annotation in Figure 1.
+Section 4.1 explains how we can compare the changes these modifications induce on the boundaries relative to rpart and PPtree using the Shiny app.
 3.1 Modification 1: Subsetting classes to produce better boundaries
-The first modification targets the third step of the original algorithm.
-Rather than combining all remaining classes into a single super-class, only the two closest ones—identified by the smallest difference in their projected means—are used to compute the next projection and define the split point
-$3^{*}$ Find an optimal one-dimensional projection $\alpha^{**}$, using a subset of $\{(\mathbf{x_{i}},y_{i}^{*})\}_{i=1}^{n}$ to separate the two class problem $g_{1}^{*}$ and $g_{2}^{*}$.
+The first modification targets the re-grouping step of the original algorithm.
+Rather than combining all remaining classes into a single super-class, only the two closest ones, as identified by the smallest difference in their projected means, are used to compute the next projection and define the split point.
+This step is now as follows:
+  • Step 2: Binary Relabeling.
+Find an optimal one-dimensional projection $\alpha^{**}$, using a subset of $\{(\mathbf{x_{i}},y_{i}^{*})\}_{i=1}^{n}$ to separate the two class problem $g_{1}^{*}$ and $g_{2}^{*}$.
 To this second projection only information from the two closest groups, one from $g_{1}^{*}$ and the other from $g_{2}^{*}$ are used.
 The best separation of $g_{1}^{*}$ and $g_{2}^{*}$ is determined by $\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{g1^{*}}<c$; then $g_{1}^{*}$ is assigned to the left node else assign $g_{2}^{*}$ to the right node, where $\mathbf{\bar{x}}_{g1^{*}}$ is the mean of $g_{1}^{*}$.
-Figure 4 illustrates the modified version of the algorithm for a three-class classification problem.
-After identifying the initial optimal projection, instead of combining two groups into a single super-group, the group furthest from the others is temporarily excluded.
-In this example, class $g_{3}$ is dropped from the current node.
-A new projection is then computed using only classes $g_{1}$ and $g_{2}$ (relabeled as $g_{1}^{*}$ and $g_{2}^{*}$ ), and one of the eight available rules is applied to determine the optimal split point.
-Subsequently, class $g_{3}$ is reintroduced, and a new projection is computed to best separate it from $g_{2}$, completing the tree construction.
-Figure 6 shows the effect of this modification on the decision boundaries obtained from the simulated dataset presented in Figure 2.
-Figure 4: Illustration of the algorithm with Modification 1 applied to a three-class problem.
-At each node, the second one-dimensional projection is computed using only the two closest groups to determine the best projection direction and split point.
+In Figure 1 this would correspond to dropping class $g_{3}$ from the current node.
+The projection is then computed using only classes $g_{1}$ and $g_{2}$ (relabeled as $g_{1}^{*}$ and $g_{2}^{*}$ ).
+The choice of the existing eight available rules is then applied to determine the optimal split point, to create the terminal node for $g_{1}$.
+Subsequently, class $g_{3}$ is reintroduced, and a new projection is computed to best separate it from $g_{2}$.
+The split is made on this set of observations, creating two more terminal nodes, completing the tree construction.
 3.2 Modification 2: Multiple splits per class based on entropy reduction
-This modification introduces a new approach for selecting the split value, based on the impurity of the resulting groups.
-This approach resembles the impurity-based criteria used in the rpart algorithm, but it operates on projected data rather than individual features.
-Currently, the only implemented impurity measure is negative entropy:
+To allow for multiple splits, various changes are needed: the split is made on the optimal projection, not after groups are combined, and new stopping rules are defined to prevent over-fitting.
+Figure 4 illustrates the change to the algorithm for this modification.
+In PPtree, the decision boundary split is made using the means for the two groups, just like is done with linear discriminant analysis.
+This modification changes this to follow the rpart algorithm, minimizing entropy, with the difference being that it is done on the optimal 1-D projected data.
+That is, the modification computes the negative entropy criteria:
 $E(s)=-\sum_{j=1}^{G}p_{js}\log(p_{js})$ (1)
 where $p_{js}$ is the proportion of observations from class $j$ in subset $s$, and $G$ is the number of classes.
-Higher values of $E(s)$ indicate greater class impurity.
-When all observations in the subset belong to the same class, the entropy reaches its minimum value of $E(s)=0$, indicating a pure node.
-To evaluate the quality of a candidate split, the impurity values of the resulting left and right subsets are combined using a weighted average of the entropie:
+Lower values of $E(s)$ indicate purity, reaching zero when all observations in the subset belong to the same class.
+The impurity value for a split is calculated by combining the values of the left and right subsets:
 $E(s_{L},s_{R})=\frac{n_{L}}{n_{L}+n_{R}}E(s_{L})+\frac{n_{R}}{n_{L}+n_{R}}E(s_{R})$ (2)
 where $n_{L},n_{R}$ are the number of observations in the left and right subsets, respectively.
-The combined impurity, $E(s_{L},s_{R})$ is computed for all possible split points-defined as the midpoints between consecutive ordered projected values-denoted $\mathscr{C}=\{c_{1},...,c_{n_{L}+n_{R}-1}\}$.
+Across all possible splits, the lowest combined impurity, $E(s_{L},s_{R})$, is computed for all possible split points, defined as the midpoints between consecutive ordered projected values, denoted $\mathscr{C}=\{c_{1},...,c_{n_{L}+n_{R}-1}\}$.
 The split point that minimizes the combined entropy is selected:
 ${\min_{c\in\mathscr{C}}}E_{c}(s_{L},s_{R})$
-This represents a major departure from the original PPtree structure, aiming to enhance flexibility in modeling non-linear class boundaries by allowing multiple splits within a class.
-Instead of relying on summary statistics at the class level, this approach shifts the focus to local differences between individual observations.
-To support this flexibility, additional control mechanisms are required, particularly new stopping criteria.
-While this is conceptually similar to the rpart algorithm, a key distinction is that larger sample sizes per node are needed to reliably estimate projection coefficients.
-Figure 5 illustrates this modified version of the algorithm, which follows the steps outlined below.
-PPtree Algorithm with Entropy-Based Splitting
+In addition, three stopping rules are introduced, which are similar to rpart: the node is pure, there are fewer than a specified number of observations in the node, or the reduction in impurity is less than a specified tolerance value.
+These changes provide the modified algorithm to be:
 Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
 Output: A projection pursuit classification tree with entropy-based node splitting
-  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates the classes in the current node.
+  • Step 1: Class Separation. Optimize a PP index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates the classes in the current node.
   • Step 2: Projection and Candidate Splits. On the projected data, $z_{i}=\boldsymbol{\alpha}^{T}\mathbf{x}_{i}$, define candidate splits as midpoints between consecutive $z_{i}$ values.
@@
 An impurity-based criterion, such as entropy, is used to determine when and how often to split observations within a node.
-4 Algorithm comparison
-To evaluate both the original algorithm and the proposed modifications, we developed an interactive Shiny app (Chang et al.
-2025).
-The tool includes methods for simulating 2D classification problems, and displays the decision boundaries generated by different algorithms: CART, PPtree, and the proposed extensions (implemented in rpart, PPtreeViz, and PPtreeExt, respectively).
-The Shiny app is available at https://natyds.shinyapps.io/explorapp/ and is implemented in PPtreeExt.
-The app contains three tabs, each tab offers control over data generation and algorithm parameters: Basic-Sim, Sim-Outlier, and MixSim.
-The first tab simulates well-separated classes, the second includes a cluster of outliers, and the third generates overlapping groups using the MixSim package (Melnykov, Chen, and Maitra 2012).
-Each tab offers control over data generation and algorithm parameters.
-The decision boundaries for rpart, PPtree, and the modified PPtree are displayed side by side for visual comparison.
-Additional controls, such as the rule partition used in PPtree algorithm and the modified PPtree based on subsetting clases, can be incorporated into the application.
-These features allow users to examine how changes in rule criteria influence the resulting classification boundaries.
-Figure 6 illustrates the use of the interactive application.
-The first modification (Subsetting clases) is applied to generate the projection, and Rule 1 is used to compute the split point.
+4 Comparative performance
+Three approaches are used to examine the effectiveness of the modifications.
+The first is a Shiny app which generates structures in 2-D and shows the model fits for three algorithms: rpart, PPtree, and the PPtree extensions.
+It was developed and used to decide on effective changes to the algorithm.
+The second is a simulation study on 94 benchmark data sets to compare test errors.
+Lastly, to understand where, why, and how the extensions work, we examine the data visually in high dimensions along with the fitted models.
+4.1 Exploring the algorithm changes interactively
+The Shiny app includes methods for simulating 2D classification problems, and displays the decision boundaries generated by three different algorithms: CART, PPtree, and the proposed extensions as implemented in the rpart, PPtreeViz, and PPtreeExt packages, respectively.
+There are three data generation methods, “Basic-Sim”, “Sim-Outlier”, and “MixSim”.
+Basic-Sim simulates classes with separations between clusters on linear combinations of variables, by user-specified correlations between variables for each cluster.
+Sim-Outlier uses the same approach with the addition of a cluster of outliers corresponding to one of the classes.
+These two methods allow only for three-cluster data.
+The user can control the number of observations in each class, the class means, and the Sim-Outlier mean and shape of the outlier groups.
+MixSim uses a Gaussian mixture model using the MixSim package (Melnykov et al.
+2012) to simulate class clusters, with any number of classes.
+While all classes have equal sample size, the user can choose this and also specify values for the MixSim algorithm.
+Figure 5, Figure 6, Figure 7 show the interface for the three different data-generating methods.
+Clicking “Ok” will simulate new samples.
+The decision boundaries for rpart, PPtree, and the modified PPtree (implemented in PPtreeExt) are displayed side by side for visual comparison.
+The user can choose which modifications of PPtree to show.
+For the other algorithms, only the default parameter choices are used.
+Figure 5: Comparing the fits for three-class data with a strong correlation between variables.
+Modification 1 (different subsetting rules) of PPtree is shown.
+The modification produces lower error and decision boundaries that are more symmetrically placed between the class clusters.
+In Figure 5, three-class data with a strong correlation between variables is simulated.
+Modification 1 (different subsetting rules) is applied, and rule 1 (Table LABEL:tbl-rules) is used to compute the split point.
 The modified PPtree achieves an error rate comparable to the original PPtree, with both substantially outperforming rpart.
-Additionally, the resulting decision boundary more accurately partitions the groups in a balanced manner.
-Figure 6: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree based on subsetting clases applied to simulated well-separated classes.
-For PPtree, Rule 1 is used to determine the split points.
-The modified PPtree produces decision boundaries that are more symmetrically placed between the clusters.
-Figure 7 presents a scenario in which Modification 2 (Multiple splits) yields a substantial improvement over the other algorithms.
-The use of multiple splits enables the small group of outliers from class 2 to be correctly classified.
-Figure 7: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree—applied to data in which one class consists of two distinct clusters.
-The boundaries produced by the modified PPtree with multiple splits better capture the underlying structure of the data compared to the other methods.
-Figure 8 compares the performance of the algorithm variants on data simulated from Gaussian mixtures.
-The user can adjust several parameters of the mixture model: $n$ is the sample size, $BarOmega$ is the desired average overlap between components, $MaxOmega$ specifies the maximum allowable overlap, and $K$ is the number of groups.
-Figure 8: Decision boundaries produced by the three algorithms— rpart, PPtree, and the modified PPtree—on data simulated from a Gaussian mixture model.
+The change in boundaries is subtle, but more satisfying.
+Between classes S (violet squares) and T (orange triangles), the boundary better matches the separation - it is not pulled away from this split by using LDA on the super-group (T and C).
+Figure 6: Comparing the fits for data where one class has two distinct clusters.
+Modification 2 (multiple splits) of PPtree is used.
+The modification produces lower error and better decision boundaries.
+Figure 6 shows data where one class has two separated clusters, as simulated by Sim-Outlier.
+The PPtree extension used is modification 2 (multiple splits).
+It yields a substantial improvement over the other algorithms, both in lower error and more sensible boundaries.
+The use of multiple splits enables the small group of outliers from class T (orange triangles) to be better classified.
+The data in Figure 7 has 5 clusters of different shapes.
+Modification 2 (multiple splits) is examined.
+Again, we see that the modified algorithm outperforms PPtree and rpart, in both lower error and better boundaries.
 The interactive application provides insight into how modifications to the algorithm affect model fit.
-It follows the principle of visualizing the model in the data space (Wickham, Cook, and Hofmann 2015).
-An additional and important use of the tool has been for debugging and evaluating the algorithmic behavior.
-By adjusting the simulated data and model parameters, it becomes evident that small changes in the data can result in substantial changes in the fitted model—particularly when projections are involved.
-This sensitivity is likely to be even greater in higher-dimensional settings.
-However, in scenarios where class separation occurs along combinations of variables, the use of projected data at each node tends to yield decision boundaries that more closely align with the true data distribution.
-5 Performance comparison
-This section presents a benchmark data study to evaluate the predictive performance of the proposed method.
-We compare the extended PPtree models with several classification methods—CART, PPtree, Random Forest, PPforest, and SVM—using 94 benchmark datasets.
-Table LABEL:tbl-source-tab1 in the Appendix contain the url with the source data while Table LABEL:tbl-data-tab2 summarize the main characteristics of each dataset summarizing the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors.
-Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes, biggest proportion size difference between classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
-For each dataset, two-thirds of the observations are randomly selected for training, and the remaining one-third is used for testing and computing the prediction error.
-This process is repeated 200 times.
-Figure 9: Average test error from 200 training/test splits of benchmark data for four models: rpart, PPtree PPtreeExt modifications 1 and 2, where the PPtree extension performs better than rpart and the original PPtree.
-Moreover, in 18 datasets, the modified PPtree outperforms CART, as illustrated in Figure 9, and in only 3 cases it achieves the lowest error among all selected methods, including ensemble-based classifiers.
-Table LABEL:tbl-datatbl presents an overview of these datasets, which is a subset of Table LABEL:tbl-data-tab2 from the Appendix.
+Comparing the fitted models follows the principle of visualizing the model in the data space (Wickham et al.
+2015).
+This app has been a valuable tool for debugging and evaluating the algorithmic behavior.
+By adjusting the simulated data and model parameters, it becomes evident that small changes in the data can result in substantial changes in the fitted model, particularly when projections are involved.
+By understanding the effect in 2-D, we aim to better understand how the algorithm might also behave in high-dimensional settings.
+Figure 7: Comparing the fits for five-cluster Gaussian mixtures. Modification 2 of PPtree is used. This produces lower error and reasonable decision boundaries.
+4.2 Numerical performance
+We compare the predictive performance for the extended PPtree algorithm with several classification methods, CART, PPtree (LDA, PDA), random forest, PPforest, and SVM, using 94 benchmark data sets.
+(LDA indicates PP conducted with the linear discriminant index, and PDA indicates PP conducted using the penalized linear discriminant index.) Figure 8 summarizes the performance on data sets where the modified algorithm outperformed other tree methods, and ?@tbl-data-tab2 summarizes the main characteristics of each data set, including the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors.
+Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes, the biggest proportion size difference between classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
+(The Appendix contains the links to the original data sources, full performance, and data set summaries.)
+In the simulation, two-thirds of the observations in each data set are randomly selected for training (using stratified sampling), and the remaining one-third is used for testing and computing the prediction error.
+This process is repeated 200 times to assess the variance in performance.
+Figure 8: Average test error with 95% error bars from 200 training/test splits of benchmark data for four models: rpart, PPtree, PPtreeExt modifications 1 and 2, where the latter performs better than the others.
+The order of the plots is based on the pattern of errors: at top left PPtreeExt (mod 2) is best but only slightly better than rpart, and at bottom right each of the PPtree versions iteratively performs better than rpart.
+Figure 8 summarises the performance for the 18 data sets, where PPtreeExt with modification 2 outperforms the other tree algorithms.
+In three of these data sets, it also has the lowest error than the ensemble-based classifiers, also.
+The plots are ordered according to the pattern of test error.
+In data sets data20-3, data27-a11, data27-a31, data27-b31, data27-c11, data27-c31, data37-5, data38, and data63, only PPtreeExt (mod 2) outperforms rpart.
+For data34 and data49-2 results are mixed.
+For the other data sets (data49-3, data22-1, data49-4, data56-1, data67, data69-1, data69-2), each of the methods does better than rpart.
+The variability across training and test splits is high for some data sets (data38, data22-1).
+Data sets data20-3, data49-2, and data69-1, representing a range of numerical performance, are chosen to investigate why PPtreeExt performs better, using high-dimensional visualisation in the next section.
 Table 2: Overview of benchmark datasets where the proposed tree-based extensions outperform the selected tree methods.
@@
 data69-2 5000 40 3 0.01 0.09
-6 Explanation of performance
-We would expect that the modified PPtree algorithm outperforms
-  • rpart when the best separation between groups is in oblique directions
-  • and the original PPtree when class variances are different, or when there are multiple classes.
-This can be investigated visually using a tour. There are several strategies for approaching this:
+4.3 Visual explanations of performance
+Exploring high dimensions is tricky, so it is helpful to set some objectives before leaping into the vast space. We would expect the modified PPtree algorithm to
+  • outperform rpart when the best separation between groups is in oblique directions
+  • and outperform PPtree when class variances are different, or when there are multiple clusters for any class.
+This can be investigated visually using a tour. There are several strategies to use for approaching this:
   1. 1.
@@
   2. 2.
-    Predict a grid of values that covers the data domain, for the methods interested in comparing, PPtreeExt, rpart and PPtree
+    Predict a grid of values that covers the data domain, for the methods interested in comparing, PPtreeExt, rpart, and PPtree, as can be done using the classifly package (Wickham 2022), and then
     1. a.
@@
     2. b.
-      record a sequence of projection bases that shows the boundary created by PPtree
+      record a sequence of projection bases that shows the boundary created by the method in focus.
     3. c.
-      use the same projection sequence to view the boundaries for the three different methods, possibly use a slice tour to focus on the boundary near the center of the data domain.
-Figure 10 show projections of the data20-3 illustrating why PPtreeExt performs better.
-Group 0 is split on either side of group 1, which means that the original PPtree algorithm would not be able to capture both clusters.
-We can also see that the separation is mostly on V3 but small amounts of other variables help to improve the separation.
-This means that rpart will not work quite as well because it will focus on splitting V3.
+      use the same projection sequence to view the boundaries for the three different methods, and possibly use a slice tour to focus on the boundary near the center of the data domain.
+Figure 9 shows two projections of the data20-3 illustrating why PPtreeExt performs better.
+The projections are 2-D from the full 5-D space.
+Color indicates the two classes (0 is blue and 1 is red).
+The circles represent the projection coefficients.
+Using a clock dial analogy, in the first projection (a), V3 is pointing towards 1 o’clock, and V4 is pointing towards 10.
+V2 is also pointing to 1, but is very small.
+The difference in the two groups is in the direction of 1, so V3 is likely mostly contributing to this, with a small contribution from V2.
+In the second projection (b) V3 points to 5, V4 points to 2, V5 points to 3, and V2 points to 9.
+The separation is in the direction of 5 o’clock, suggesting again that V3 is primarily important.
+Group 0 is split on either side of group 1, which means that the PPtree algorithm would not be able to capture both clusters, but modification 2 would.
+The separation between the two classes is mostly on V3 but small contributions from other variables help to improve the separation.
+This means that rpart will not work quite as well because it will focus on splitting V3, ignoring the other variables.
 (a) projection 1
@@
 Figure 9: Two projections from a grand tour of data20-3, from which we can see why the PPtreeExt method is the most effective.
-The group 0 is split into two clusters on either side of group 1, which means that the original algorithm is not optimal.
-Secondly, the separation between the two groups is in a combination of variables.
-Figure 11 shows some plots illustrating why PPtreeExt outperforms the other methods of the data49-2 data.
-From Figure 9 we can see that the original PPtree (with LDA) algorithm performs much better than a tree, and the PPtreeExt performs much better again.
-This data has 100 predictors and 2 classes (Table LABEL:tbl-datatbl), which makes visualising more challenging that data20-3.
+Class 0 is split into two clusters on either side of group 1, which means that the original algorithm is not optimal.
+Secondly, the separation between the two groups is primarily V3 with small contributions from other variables, which gives PPtreeExt a slight edge over rpart.
+From Figure 8, for data49-2, we can see that the original PPtree (with LDA) algorithm performs much better than a tree, and the PPtreeExt performs much better again.
+This data has 100 predictors and 2 classes (Table LABEL:tbl-datatbl), which makes visualising more challenging than data20-3.
 A useful approach is to pre-process using principal component analysis (PCA) on the predictors, to reduce the number of variables to visualise, while maintaining class structure.
-Plot (a) shows PC 1 vs PC2, with points coloured by the two classes.
-PC 1 explains 99.8% of the variability in the predictors but it explains nothing of the class structure.
-This is not uncommon for PCA, where often interesting structure can be found in the smaller components.
+In Figure 10 plot (a) shows PC 1 vs PC2, with points coloured by the two classes.
+PC 1 explains 99.8% of the variability in the predictors, but it explains nothing of the class structure.
+This is not uncommon for PCA, where an interesting structure can be found in the smaller components.
 Plot (b) shows PC 2 vs PC 3, and here we see interesting class differences.
 Each class appears to have three linear pieces pointing in different directions.
-This supports why PPtreeExt might work well, because in the combination of variables, the classes are nearly separable with multiple linear cuts.
-It’s not feasible to continue working through principal components to tease apart the cluster structure relating to classes, so a good next step is to do projection pursuit.
+This supports why PPtreeExt might work well: in the combination of variables, the classes are nearly separable with multiple linear cuts.
+It’s not feasible to continue working through principal components to tease apart the cluster structure relating to classes, so a good next step is to do PP with the guided tour.
 The optimal projection is shown in plot (c).
-It can be seen the two classes are mostly separable by a linear combination of variables, but there is no gap between the clusters.
+It is computed on 26 PCs, and many of these contribute to this projection as seen by the mess of lines inside the circle representing the axes.
+The two classes are mostly linearly separable in this projection, but there is no clear gap between the clusters.
 The linear combination with so many variables is what causes some difficulties for a simple tree algorithm.
-The interesting shapes of the clusters for each class, which appear to be multiple linear pieces, means that the PPtreeExt can find better separations than PPtree.
+The interesting shapes of the clusters for each class, which appear to be multiple linear pieces, mean that the PPtreeExt can find better separations than PPtree.
 (a) PC1 vs PC2
@@
 PC1 explains 99.8% of the overall variability among predictors, but it is not useful for separating the classes (a).
-PC2 and PC3 (b) have an interesting difference between the two classes which shows why multiple linear splits might be useful.
-A guided projection pursuit on the first 26 principal components (c) shows that these classes could be separated with a linear combination.
+PC2 and PC3 (b) have an interesting difference between the two classes, which shows why multiple linear splits might be useful.
+A guided projection pursuit tour on the first 26 principal components (c) shows that these classes could be separated with a linear combination.
 This is why the original PPtree classifier does better than the tree algorithm and also why the extension improves the performance.
-Figure 12 has plots illustrating why the PPtreeExt performs better on data69-1.
-From Figure 9 we see that all of the PPtree versions perform better than an rpart fit.
+From Figure 8 we see that all of the PPtree versions perform better than an rpart fit on data69-1.
 This data has 21 predictors and 3 classes (Table LABEL:tbl-datatbl).
-This is sufficiently small to examine using a tour, and we initially use a guided tour with the LDA index.
-This produced the projection shown in plot (a) where we see the primary difference between the three classes.
-These are each similar size, but oriented in such a way that in a 2D projection they form a triangle.
+This is a sufficiently small-scale problem to examine using a tour, and we initially use a guided tour with the LDA index.
+This produced the projection shown in plot (a), where we see the primary difference between the three classes.
+These are each similar size, but oriented in such a way that in this 2-D projection, they form a triangle.
 Plots (b) and (c) show the boundaries using a slice tour with the same projection as used in (a) for PPtreeExt and rpart, respectively.
-The boundary points are generated by simulating a random sample of observations between (-3, 3) in each dimension.
-The slice cuts through to centre of the box.
-We can see that PPtreeExt essentially fits the data shape neatly but the rpart boundary is very messy.
-This data needs oblique cuts, which PPtreeExt provides, and the modified algorithm handles the variance differences better than the original.
+This uses approach 2 above: created by predicting the class of points randomly generated from a uniform distribution in a 21-D cube of the same range as the data.
+The slice cuts through to the center of the box.
+We can see that PPtreeExt essentially fits the data shape neatly, but the rpart boundary is very messy.
+This data needs oblique cuts, which PPtreeExt provides but rpart does now.
+The modified algorithm handles the variance differences better than the original PPtree.
 (a) data
@@
 (c) rpart
-Figure 12: Examining why data49-2 is best classified using PPtreeExt: (a) best projection from a guided tour shows that the data has three similar linear pieces oriented to form a triangle, (b) boundary produced by the PPtreeExt fitted model and (c) boundary produced by rpart fitted model.
+Figure 11: Examining why data69-1 is best classified using PPtreeExt: (a) best projection from a guided tour shows that the data has three similar linear pieces oriented to form a triangle, (b) boundary produced by the PPtreeExt (mod 2) fitted model, and (c) boundary produced by rpart fitted model.
+The same projection is used in all plots to allow comparison.
+The boundary is examined by predicting the class of a uniform random sample of points in the data space, and viewed using a slice tour.
+The boundary for PPtreeExt (mod 2) mirrors what we learn from the data, that it is carving the data space into three equiangular chunks, but rpart is not capturing these linear combinations.
 5 Discussion
-This article has presented two modifications to the PPtree algorithm for classification problems implemented in the R package PPtreeExt and evaluated using a purpose-built interactive web application and real data.
-The first modification involves subsetting classes when calculating the projection direction at each node, while the second allows for multiple splits per class based on individual entropy evaluations between candidate splits.
+This article has presented two modifications to the PPtree algorithm for classification problems implemented in the R package PPtreeExt.
 These changes result in a more flexible classification method that leverages linear combinations of variables.
-The web application also demonstrates how easily diagnostic tools can be constructed to explore and understand the behavior of new algorithms.
-The benchmark data study showed that PPtreeExt predictive performance is better than, CART and PPtree in 18 data out of 94.
-The modified PPtree algorithm outperforms CART in cases when the best separation between groups is in oblique directions and outperforms the original PPtree when class variances are different, or when there are multiple classes.
-Interpretability is increasingly important, because predictions from algorithms may need to be defended in practice.
-We have shown how to examine the data and the fitted model to understand why the PPtree extension performs better than other tree methods on some data sets.
+The performance study showed that the predictive performance of PPtreeExt with modification 2 is better than CART and PPtree in 18 out of 94 data sets.
+A web app was developed and shared to help understand and debug modifications to the algorithm.
+Interpretability is increasingly important because predictions from algorithms may need to be defended in practice.
+We have shown how to examine the data and the fitted model using high-dimensional visual methods to understand why the PPtree extension performs better than other tree methods on some data sets.
 This approach is generally useful for assessing and diagnosing model fits.
-The modified tree algorithm can be incorporated into a projection pursuit forest classifier (da Silva, Cook, and Lee 2025).
+Ultimately, this may provide better building blocks for a projection pursuit-based classification random forest algorithm (da Silva et al.
+2021) as available in the PPforest R package (da Silva et al.
+2025).
+The modified tree algorithm can be incorporated into a projection pursuit forest classifier (da Silva et al.
+2025).
 Given its implicit variable selection mechanism, the ensemble is expected to handle nuisance variables more effectively than a single tree.
+Using the PDA index (penalised discriminant analysis) also helps when there are many variables.
 Additionally, the use of bagging should yield improved and more robust performance in problems with nonlinear decision boundaries.
 This can be accomplished using the current PPforest package.
-8 Appendix
-Table 3: Source data information
-Name URL
-data12 https://archive.ics.uci.edu/dataset/518/speaker+accent+recognition
-data13-1 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
-data13-2 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
-data20-1 https://archive.ics.uci.edu/dataset/357/occupancy+detection
-data20-2 https://archive.ics.uci.edu/dataset/357/occupancy+detection
-data20-3 https://archive.ics.uci.edu/dataset/357/occupancy+detection
-data21 https://archive.ics.uci.edu/dataset/537/cervical+cancer+behavior+risk
-data22-1 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
-data22-2 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
-data26 https://archive.ics.uci.edu/dataset/308/gas+sensor+array+under+flow+modulation
-data27-a11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-a21 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-a31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-a32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-b11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-b12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-b31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-b32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-c11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-c12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-c31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data27-c32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
-data29 https://archive.ics.uci.edu/dataset/282/lsvt+voice+rehabilitation
-data32 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
-data34 https://archive.ics.uci.edu/dataset/267/banknote+authentication
-data36 https://archive.ics.uci.edu/dataset/236/seeds
-data37-1 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
-data37-2 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
-data37-3 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
-data37-4 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
-data37-5 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
-data37-8 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
-data37-14 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
-data38 https://archive.ics.uci.edu/dataset/230/planning+relax
-data39-1 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
-Name URL
-data39-2 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
-data39-3 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
-data39-4 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
-data39-5 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
-data39-6 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
-data39-7 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
-data39-8 https://archive.ics.uci.edu/dataset/225/gas+sensor+array+drift+dataset
-data39-9 https://archive.ics.uci.edu/dataset/226/gas+sensor+array+drift+dataset
-data41-1 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
-data41-2 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
-data41-3 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
-data42 https://archive.ics.uci.edu/dataset/192/breast+tissue
-data44 https://archive.ics.uci.edu/dataset/181/libras+movement
-data49-2 https://archive.ics.uci.edu/dataset/167/hill+valley
-data49-3 https://archive.ics.uci.edu/dataset/168/hill+valley
-data49-4 https://archive.ics.uci.edu/dataset/169/hill+valley
-data50 https://archive.ics.uci.edu/dataset/161/mammographic+mass
-data52-1 https://archive.ics.uci.edu/dataset/96/spectf+heart
-data52-2 https://archive.ics.uci.edu/dataset/96/spectf+heart
-data53 https://archive.ics.uci.edu/dataset/139/synthetic+control+chart+time+series
-data55-1 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
-data55-2 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
-data56-1 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
-data56-2 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
-data57 https://archive.ics.uci.edu/dataset/39/ecoli
-data61-1 https://archive.ics.uci.edu/dataset/54/isolet
-data62 https://archive.ics.uci.edu/dataset/74/musk+version+1
-data63 https://archive.ics.uci.edu/dataset/75/musk+version+2
-data64-1 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
-data64-2 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
-data65 https://archive.ics.uci.edu/dataset/15/breast+cancer+wisconsin+original
-data66 https://archive.ics.uci.edu/dataset/62/lung+cancer
-data67 https://archive.ics.uci.edu/dataset/59/letter+recognition
-data68-1 https://archive.ics.uci.edu/dataset/50/image+segmentation
-data68-2 https://archive.ics.uci.edu/dataset/50/image+segmentation
-data69-1 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
-data69-2 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
-data70 https://archive.ics.uci.edu/dataset/42/glass+identification
-data71-1 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
-Name URL
-data71-2 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
-data72-1 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
-data72-2 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
-data72-3 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
-data72-4 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
-data72-5 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
-data72-6 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
-data72-7 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
-data73 https://archive.ics.uci.edu/dataset/151/connectionist+bench+sonar+mines+vs+rocks
-data74 https://archive.ics.uci.edu/dataset/152/connectionist+bench+vowel+recognition+deterding+data
-NCI60 https://github.com/natydasilva/PPforest/blob/master/data/NCI60.rda
-crab https://github.com/natydasilva/PPforest/blob/master/data/crab.rda
-fishcatch https://github.com/natydasilva/PPforest/blob/master/data/fishcatch.rda
-glass https://github.com/natydasilva/PPforest/blob/master/data/glass.rda
-image https://github.com/natydasilva/PPforest/blob/master/data/image.rda
-leukemia https://github.com/natydasilva/PPforest/blob/master/data/image.rda
-lymphoma https://github.com/natydasilva/PPforest/blob/master/data/lymphoma.rda
-olive https://github.com/natydasilva/PPforest/blob/master/data/olive.rda
-parkinson https://github.com/natydasilva/PPforest/blob/master/data/parkinson.rda
-wine https://github.com/natydasilva/PPforest/blob/master/data/wine.rda
-Table 6: Overview of benchmark datasets
-dataset Cases Predictors Groups Imbalance Correlation
-data12 329 12 6 0.41 0.38
-data13-1 122 13 2 0.03 0.41
-data13-2 121 13 2 0.29 0.37
-data20-1 2665 5 2 0.27 0.81
-data20-2 9752 5 2 0.58 0.29
-data20-3 8143 5 2 0.58 0.45
-data21 72 19 2 0.42 0.22
-data22-1 258 5 4 0.25 0.12
-data22-2 145 5 4 0.14 0.18
-data26 58 435 4 0.21 0.64
-data27-a11 1747 18 5 0.38 0.24
-data27-a21 1264 18 5 0.35 0.22
-data27-a31 1834 18 5 0.28 0.27
-data27-a32 1830 32 5 0.28 0.13
-data27-b11 1072 18 5 0.31 0.27
-data27-b12 1069 32 5 0.32 0.13
-data27-b31 1424 18 5 0.21 0.24
-data27-b32 1420 32 5 0.21 0.11
-data27-c11 1111 18 5 0.14 0.27
-data27-c12 1107 32 5 0.13 0.11
-data27-c31 1448 18 5 0.17 0.27
-data27-c32 1444 32 5 0.17 0.11
-data29 126 310 2 0.33 0.28
-data32 403 6 4 0.20 0.11
-data34 1372 4 2 0.11 0.43
-data36 210 7 3 0.00 0.61
-data37-1 29101 53 5 0.23 0.17
-data37-14 769 53 2 0.52 0.16
-data37-2 14132 53 3 0.62 0.18
-data37-3 11835 53 4 0.34 0.16
-data37-4 16357 53 5 0.29 0.15
-data37-5 17560 53 5 0.21 0.16
-dataset Cases Predictors Groups Imbalance Correlation
-data37-8 22990 53 9 0.27 0.17
-data38 182 12 2 0.43 0.21
-data39-1 445 128 6 0.15 0.55
-data39-2 1239 128 5 0.35 0.40
-data39-3 1586 128 5 0.17 0.56
-data39-4 161 128 5 0.32 0.43
-data39-5 197 128 5 0.22 0.57
-data39-6 2300 128 6 0.25 0.61
-data39-7 3613 128 6 0.11 0.68
-data39-8 294 128 6 0.43 0.57
-data39-9 470 128 6 0.10 0.42
-data41-1 5456 2 4 0.34 0.03
-data41-2 5456 4 4 0.34 0.15
-data41-3 5456 24 4 0.34 0.17
-data42 106 9 6 0.08 0.51
-data44 360 90 15 0.00 0.27
-data49-2 606 100 2 0.03 1.00
-data49-3 606 100 2 0.01 0.99
-data49-4 606 100 2 0.01 0.99
-data50 830 5 2 0.03 0.23
-data52-1 80 44 2 0.00 0.25
-data52-2 187 44 2 0.84 0.29
-data53 600 60 6 0.00 0.46
-data55-1 3823 64 10 0.00 0.12
-data55-2 1797 64 10 0.01 0.12
-data56-1 7494 16 10 0.01 0.27
-data56-2 3498 16 10 0.01 0.28
-data57 327 7 5 0.38 0.23
-data61-1 6238 617 26 0.00 0.17
-data62 476 166 2 0.13 0.26
-data63 6598 166 2 0.69 0.27
-data64-1 4435 36 6 0.15 0.49
-data64-2 2000 36 6 0.13 0.50
-data65 683 9 2 0.30 0.60
-data66 27 56 3 0.07 0.20
-dataset Cases Predictors Groups Imbalance Correlation
-data67 20000 16 26 0.00 0.18
-data68-1 210 19 7 0.00 0.30
-data68-2 2100 19 7 0.00 0.28
-data69-1 5000 21 3 0.01 0.30
-data69-2 5000 40 3 0.01 0.09
-data70 214 9 6 0.31 0.23
-data71-1 43500 9 7 0.78 0.19
-data71-2 14494 9 5 0.79 0.20
-data72-1 94 18 4 0.09 0.39
-data72-2 94 18 4 0.05 0.40
-data72-3 94 18 4 0.10 0.42
-data72-4 94 18 4 0.12 0.42
-data72-5 94 18 4 0.10 0.49
-data72-6 94 18 4 0.06 0.40
-data72-7 94 18 4 0.06 0.44
-data73 208 60 2 0.07 0.23
-data74 990 12 11 0.00 0.22
-NCI60 56 31 7 0.05 0.56
-crab 200 6 4 0.00 0.95
-fishcatch 159 7 7 0.31 0.46
-glass 214 10 6 0.31 0.23
-image 2310 19 7 0.00 0.28
-leukemia 72 41 3 0.40 0.44
-lymphoma 80 51 3 0.41 0.75
-olive 572 9 9 0.32 0.35
-parkinson 195 23 2 0.51 0.50
-wine 178 14 3 0.13 0.30
+6 Supplementary Materials
+This article was written in Quarto (Allaire et al.
+2024), using the R package ggplot2 (Wickham 2016) for visualizations.
+The files to reproduce the article and results are available at https://github.com/natydasilva/PPtreeExt_paper_JCGS.
+Animated gifs corresponding to Figure 9, Figure 10 and Figure 11 are in the gifs folder.
+The code directory contains the scripts for the simulations and to explore the model fits in high dimensions.
+The PPtreeExt package provides the modified PPtree algorithm.
+The Shiny app is available at https://natyds.shinyapps.io/explorapp/ and the app code is provided in the PPtreeExt package, so that it can be used locally for new algorithm developments.